OpenStudy (anonymous):

ok please help its review Simplify: 2x(4x2 - 3x - 9) and Multiply: 5xy(2x - y)

OpenStudy (anonymous):

@UnkleRhaukus could u help

OpenStudy (unklerhaukus):

\[2x(4x^2 - 3x - 9)\]\[=(2x\times4x^2)+(2x\times-3x)+(2x\times-9)\] \[=\cdots\]

OpenStudy (unklerhaukus):

\[5xy(2x - y)=(5xy\times2x)+(5xy\times-y)=\cdots\]

OpenStudy (anonymous):

? 1 i think the answer is 8x3 - 3x2 - 9x

OpenStudy (unklerhaukus):

almost dont forget to multiply the 2 next to the x

OpenStudy (anonymous):

i dont understand

OpenStudy (unklerhaukus):

\[=(2x\times4x^2)+(2x\times-3x)+(2x\times-9)\]\[(8)x^3+(2\times-3)x^2+(2\times-9)x\]

OpenStudy (anonymous):

: 2x(4x2 - 3x - 9) A. 8x2 - 3x - 9 B. 8x2 - 6x - 18 C. 8x3 - 6x2 - 18x D. 8x3 - 3x2 - 9x

OpenStudy (anonymous):

Multiply: 5xy(2x - y) A. 10xy - 5xy B. 10x2y - 5xy2 C. 10x2 - 5xy + 25y2 D. 10x2 - xy + y2

OpenStudy (unklerhaukus):

simplify the terms in the bracket in my most recent line

OpenStudy (anonymous):

i dont get it